OpenStudy (anonymous):

solve for X 5/x^2-1 - 2/x = 2/x+1

OpenStudy (zzr0ck3r):

can you rewrite that with more ().

hero (hero):

1. Make sure the denominators are factored: \(\frac{5}{(x+1)(x-1)} - \frac{2}{x} = \frac{2}{x+1}\) 2. Multiply both sides by x-1 to get: \(\frac{5}{x+1}- \frac{2x-2}{x} = \frac{2x-2}{x+1}\) 3. Add (2x-2)/x to both sides; subtract (2x-2)/(x+1) from both sides to get: \(\frac{5}{x+1} - \frac{2x-2}{x+1} = \frac{2x-2}{x}\) 4. Combine fractions on the left side to get \(\frac{7 - 2x}{x+1} = \frac{2(x-1)}{x}\) 5. Cross multiply to get: \(x(7-2x) = 2(x+1)(x-1)\) 6. Simplify and Multiply to get \(7x - 2x^2 = 2(x^2 -1)\) 7. Simplify further to get: \(7x-2x^2 = 2x^2 - 2\) Add 2x^2 to both sides; subtract 7x from both sides to get: \(4x^2 - 7x - 2 = 0\) Finish solving the quadratic equation.

OpenStudy (anonymous):

Thanks